# Vital Insight: Assisting Experts’ Context-Driven Sensemaking of Multi-modal Personal Tracking Data Using Visualization and Human-In-The-Loop LLM Agents

### v2

@@ 8. Experts’ sensemaking model
Experts frequently compare various sensing modalities and integrate them to synthesize patterns. We identified five types of patterns that experts pay attention to: long duration, change, discrepancy, gap, and routine activities. To explain these patterns, experts turn to contextual information like user profiles and check-ins, and use them to fill in gaps that sensing data alone cannot provide, often involving a process of extracting relevant details. Based on the patterns, experts may take certain actions directly, such as addressing sensor issues. More commonly, however, experts first work to synthesize a user persona before taking action. These personas are typically constructed from patterns observed in longitudinal patterns, along with known contextual information such as demographics. Common persona traits include lifestyle, hobbies, social interactions, and regular routines. Once the persona is established, experts use it to guide specific actions, such as reaching out to relevant stakeholders or designing personalized interventions. Furthermore, the sensemaking procedures described above can be supported by AI inference at varying levels of granularity. For instance, observation summaries of sensing data can help experts compare and integrate raw data across modalities to identify patterns, while more creative inferences can assist them in brainstorming potential explanations for those patterns.
@@
To further elevate the sensemaking model, we consider the collaboration between experts, data, and AI. Experts iteratively move between direct data representation and indirect inference to generate insights (Fig. 8). A common approach is exploration, where experts examine the data directly and develop their thoughts. Two other main paths involve indirect inference. One starts with questioning certain aspects of a indirect inference and seeking evidence from data to validate the results. The other involves finding inspiration, such as overlooked details in indirect inference, and then intentionally retrieving relevant information from the data. These three paths lead to insights, where experts feel confident in drawing conclusions. After reaching insights, some experts may revisit the indirect inference to confirm their assumptions. In addition to generating insights, these paths may also lead to confusion. Some experts choose to consult indirect inference for hints while others routed to stakeholders like participants to reach confident answers.
@@ 9.1.4. Provide Both AI Assisted Contextual Overview and Momentary Highlights
Across our studies, we found that AI assistance should provide both a semantic overview and momentary highlights to facilitate the iterative sensemaking process. The contextual overview, such as a daily summary, serves as the background context that is “always present,” almost like an additional data modality that provides an overview of the person’s day on a higher level. The momentary highlights, such as occurrences at specific times in VI-2, integrate well with other sensing data on the same time scale, and help experts quickly navigate important information. Both functionalities are essential for experts’ sensemaking process, complementing and enhancing each other, and should be integrated in a successful application.
@@
To further examine various sensemaking processes and goals, Faisal et al. categorized six types of sensemaking representations (faisal2009classification, ). In the specific context of making sense of multi-modal health sensing data, experts primarily use argumentational representations, which are formed by integrating a series of claims to infer conclusions or establish inferential relationships. Since health-sensing data related to human behavior is often time-based, many sensemaking processes fall under sequential representations. Chronologies, for example, are a common type of sequential representation. As noted in our findings, experts often examine data from beginning to end. However, non-sequential representations frequently emerge iteratively, where human-in-the-loop AI augmentation can help experts quickly observe patterns. Another category of sensemaking we observed is faceted representation, often used in abductive reasoning, where humans try to derive the most reasonable explanation from the data. The authors describe this process as finding or adding different (often new) entities that influence sensemaking (faisal2009classification, ). In our study, the entities—sensing modalities—are fixed, meaning researchers theoretically cannot acquire additional data. However, this process still applies from a different perspective: the abundance of modalities can make it challenging to become familiar with all of them at once. Experts typically start with a subset of modalities and gradually explore additional information as needed. Beyond modalities, “the place to start” also involves what we defined as “occurrences” in Study 3 – specific events that immediately draw attention. Another element that could be considered a facet is AI assistance, which can be viewed as an augmentation of data and visualization, as described by Wu et al. (wu2021ai4vis, ). In our study, particularly in Study 3, AI assistance occasionally transcends mere data augmentation, becoming a unique facet of the data source within experts’ sensemaking models. For example, after becoming familiar with the system, many experts prioritize the AI’s daily summaries or explanations of occurrences to guide their sensemaking process. They may then refer back to the raw data, rather than first examining the data and consulting AI content only when necessary. These interactions were described as the ”inspire-retrieve” route in Fig. 8. Although AI does not generate new data, it is crucial to recognize its standalone role as a new facet of information within human-in-the-loop sensemaking processes.
@@ 9.3. Biases in LLM-generated inferences
During user testing with experts, we observed them gradually building trust in the LLM-generated inferences and summaries while exploring the various system components. However, this process also increased the potential risk of bias introduced by inaccurate LLM results. Several prior works have discussed concerns with bias in LLM-generated outcomes: highlighting biases related to gender (wan2023kelly, ), age (duan2024large, ), geography (manvi2024large, ), and politics (urman2023silence, ). These biases could influence the interpretation of contextual information about individuals. For example, LLM may offer a stereotyped interpretation of the data based on the user’s certain demographics. LLMs also exhibit bias in their information retrieval processes, such as giving preference to documents or items from specific input positions like the beginning or end of the list (dai2024neural, ; li2023halueval, ; dai2024bias, ). This bias could potentially affecting the retrieval of various raw sensing data. Since our user testing primarily involved people with extensive expertise, they expressed a highly skeptical view of the LLM-generated results, often validating the inferences before accepting them. For example, P2 specifically mentioned concerns about being influenced by the LLM summary and stated they would avoid checking it first to prevent bias. However, as their trust in the system grew, there is a risk that they might be increasingly influenced by the LLM’s outputs, even for experts (choudhury2024large, ; kerasidou2022before, ; ryan2020ai, ). During our internal evaluation of the LLM-generated results, we also noted a potential concern with over-interpreting sensor data, where the LLM might notice trivial fluctuations in heart rate and force itself to make an inference. Without explicit instructions, LLMs also tend to not question the validity and reliability of the data and always offer a potential explanation regardless. Future research must recognize the importance of examining both the LLM’s biases in interpreting personal tracking data and the biases experts might develop when integrating those results.
@@ 9.4. Designing LLM agents that mimic experts’ sensemaking process
Although this study did not aim to enhance the accuracy of complex activity recognition and sensemaking, our insights into experts’ sensemaking processes may help inform the design of more effective AI-assisted systems for interpreting personal tracking data. Guiding LLMs with expert knowledge is a common approach for solving domain-specific problems, including techniques like adding instructions in prompts (ji2024hargpt, ; cosentino2024towards, ), creating expert agents (yang2024drhouse, ), and fine-tuning (kim2024health, ; cosentino2024towards, ). In the context of making sense of sensor data, HARGPT employs simple role-playing prompts without expert guidance to identify activities from raw IMU data (ji2024hargpt, ). Health-LLM evaluated zero/few-shot learning and fine-tuning with simple prompts on sensor data (kim2024health, ). Similarly, prior works have also explored how to encode sensor data to prompts: LLMSense and PhysioLLM use text-formatted sensor data and prompts to derive high-level inferences from sensor traces (ouyang2024llmsense, ; fang2024physiollm, ); Cosetino et al. incorporated domain knowledge into prompts and fine-tuned models using expert responses (cosentino2024towards, ). We build on previous works to provide an iterative and evolving approach to continuously include domain expertise as Data Interpretation Guidance in the prompt. While prior works have explored incorporating domain/expert knowledge, we observed that much of the expert knowledge is usually “static.”. From related work, it is clear that the complex sensemaking processe experts use when interpreting personal tracking data are under-documented and rarely considered to guide LLMs. Our study reveals that these sensemaking processes extend beyond basic knowledge, like the normal range of a person’s heart rate, and encompass the selection and integration of multiple data streams, handling data discrepancies and uncertainties, and strategies for information retrieval. These domain-specific insights can guide LLMs in various ways. For instance, understanding how experts initiate sensemaking – starting from changes rather than following a strict chronological order – could be used to optimize data retrieval for LLMs. Additionally, the differing definitions of anomalies between the LLM and experts highlight an additional area for improvement, involving extending the system to incorporate more task-specific and sophisticated anomaly detection algorithms. Additionally, our findings on how experts collaborate with LLMs and other informational modules to reach conclusions underscore how LLM detection models should be designed for real-world applications to support experts effectively. We hope that, although not the main focus of this study, our insights will help AI and LLM researchers recognize opportunities to enhance current detection systems and better support complex sensemaking tasks.
@@ 9.6. Limitations
The study has several limitations worth noting. While we achieved data saturation, our small sample size limits the generalizability of our findings, particularly as it relates to other health and behavioral outcomes. Although recent research suggests small sample sizes are acceptable in interpretive studies (crabtree2024h, ), future work should involve larger populations and diverse scenarios to validate the findings. While researchers reviewed all the LLM-generated content showed in the two user studies to ensure basic correctness without making any changes, we did not conduct a quantitative evaluation of the standalone LLM results. A more thorough evaluation of LLM results and capabilities can be conducted in future research, which is beyond the scope of this study. Another limitation arises from the fact that experts spent more time on tasks when LLM content was included compared to when it was not. In our qualitative analysis, we found that this was partly due to experts treating LLM content seriously in their sensemaking process, which led them to spend more time on it. Additionally, although we provided a tutorial video at the beginning, there was still a learning curve in using the information effectively. A more longitudinal study is needed to assess the long-term usage of the prototype in real-world settings and determine whether it can ultimately save experts time in sensemaking tasks.

### v3

@@ 8. Experts’ sensemaking model
Experts frequently compare various sensing modalities and integrate them to synthesize patterns. We identified five types of patterns that experts pay attention to: long duration, change, discrepancy, gap, and routine activities. To explain these patterns, experts turn to contextual information like user profiles and check-ins, and use them to fill in gaps that sensing data alone cannot provide, often involving a process of extracting relevant details. Based on the patterns, experts may take certain actions directly, such as addressing sensor issues. More commonly, however, experts work to synthesize a user persona before taking action. These personas are typically constructed from patterns observed in longitudinal patterns, along with known contextual information such as demographics. Common persona traits include lifestyle, hobbies, social interactions, and regular routines. Once the persona is established, experts use it to guide specific actions, such as reaching out to relevant stakeholders or designing personalized interventions. Furthermore, the sensemaking procedures described above can be supported by AI inference at varying levels of granularity. For instance, observation summaries of sensing data can help experts compare and integrate raw data across modalities to identify patterns, while more creative inferences can assist them in brainstorming potential explanations for those patterns.
@@
To further elevate the sensemaking model, we consider the collaboration between experts, data, and AI. Experts iteratively move between direct data representation and indirect inference to generate insights (Fig. 9). A common approach for experts is to explore the data directly and develop their thoughts. Two other main paths involve indirect inference. One starts with questioning certain aspects of an indirect inference and seeking evidence from data to validate the results. The other involves being inspired by AI-supported inference, such as overlooked details in indirect inference, and then intentionally retrieve relevant information from the data. These three paths lead to insights, where experts feel confident in drawing conclusions. After reaching insights, some experts may revisit the indirect inference to confirm their assumptions. In addition to generating insights, these paths may also lead to confusion. Some experts choose to consult indirect inference for hints while others routed to stakeholders like participants to reach confident answers.
Experts’ sensemaking model also highlights the need for customization. While there are many commonalities in their mental models and thought processes, each expert’s focus varies slightly. Differences include starting points, preferred modalities, the amount of data or time range they are comfortable reviewing, familiarity with the target population, and preferred processing modes (e.g., more visual or more text-based). Based on this model, the system should allow extensive customization across these aspects, including basic interface customization like layout and color following Schneiderman’s mantra “Overview first, zoom and filter, then details-on- demand” (shneiderman2003eyes, ), as well as advanced customization based on the personal sensemaking model. It’s also important to recognize the varying levels of trust in AI-generated content among different experts. Systems should tailor the tone of the AI to be more conservative, with additional proof and explanations for certain experts, while provide more insights from the LLM for those open to getting inspiration from the outputs.
@@ 9.1.4. Provide Both AI Assisted Contextual Overview and Momentary Highlights
Across our studies, we found that AI assistance should provide both a semantic overview and momentary highlights to facilitate the iterative sensemaking process. The contextual overview, such as a daily summary, serves as the background context that is “always present”, almost like an additional data modality that provides an overview of the person’s day on a higher level. The momentary highlights, such as occurrences at specific times in VI-2, integrate well with other sensing data on the same time scale, and help experts quickly navigate important information. Both functionalities are essential for experts’ sensemaking process, complementing and enhancing each other, and should be integrated in a successful application.
@@
To further examine various sensemaking processes and goals, Faisal et al. categorized six types of sensemaking representations (faisal2009classification, ). In the specific context of making sense of multi-modal health sensing data, experts primarily use argumentational representations, which are formed by integrating a series of claims to infer conclusions or establish inferential relationships. Since health-sensing data related to human behavior is often time-based, many sensemaking processes fall under sequential representations. Chronologies, for example, are a common type of sequential representation. As noted in our findings, experts often examine data from beginning to end. However, non-sequential representations frequently emerge iteratively, where human-in-the-loop AI augmentation can help experts quickly observe patterns. Another category of sensemaking we observed is faceted representation, often used in abductive reasoning, where humans try to derive the most reasonable explanation from the data. The authors describe this process as finding or adding different (often new) entities that influence sensemaking (faisal2009classification, ). In our study, the entities—sensing modalities—are fixed, meaning researchers theoretically cannot acquire additional data. However, this process still applies from a different perspective: the abundance of modalities can make it challenging to become familiar with all of them at once. Experts typically start with a subset of modalities and gradually explore additional information as needed. Beyond modalities, “the place to start” also involves what we defined as “occurrences” in Study 3 – specific events that immediately draw attention. Another element that could be considered a facet is AI assistance, which can be viewed as an augmentation of data and visualization, as described by Wu et al. (wu2021ai4vis, ). In our study, particularly in Study 3, AI assistance occasionally transcends mere data augmentation, becoming a unique facet of the data source within experts’ sensemaking models. For example, after becoming familiar with the system, many experts prioritize the AI’s daily summaries or explanations of occurrences to guide their sensemaking process. They may then refer back to the raw data, rather than first examining the data and consulting AI content only when necessary. These interactions were described as the “inspire-retrieve” route in Fig. 9. Although AI does not generate new data, it is crucial to recognize its standalone role as a new facet of information within human-in-the-loop sensemaking processes.
@@ 9.3. Biases in LLM-generated inferences
During user testing with experts, we observed them gradually building trust in the LLM-generated inferences and summaries while exploring the various system components. However, this process also increased the potential risk of bias introduced by inaccurate LLM results. Several prior works have discussed concerns with bias in LLM-generated outcomes: highlighting biases related to gender (wan2023kelly, ), age (duan2024large, ), geography (manvi2024large, ), and politics (urman2023silence, ). These biases could influence the interpretation of contextual information about individuals. For example, LLM may offer a stereotyped interpretation of the data based on the user’s certain demographics. LLMs also exhibit bias in their information retrieval processes, such as giving preference to documents or items from specific input positions like the beginning or end of the list (dai2024neural, ; li2023halueval, ; dai2024bias, ). This bias could potentially affecting the retrieval of various raw sensing data. Since our user testing primarily involved people with extensive expertise, they expressed a highly skeptical view of the LLM-generated results, often validating the inferences before accepting them. For example, P2 specifically mentioned concerns about being influenced by the LLM summary and stated they would avoid checking it first to prevent bias. However, as their trust in the system grew, there is a risk that they might be increasingly influenced by the LLM’s outputs, even for experts (choudhury2024large, ; kerasidou2022before, ; ryan2020ai, ). During our internal evaluation of the LLM-generated results, we also noted a potential concern with over-interpreting sensor data, where the LLM might notice trivial fluctuations in heart rate and force itself to make an inference. Without explicit instructions, LLMs also tend to not question the validity and reliability of the data and always offer a potential explanation regardless. While experts may be able to identify incorrect outputs from the LLM, there could still be a subtle bias introduced. Even without considering incorrect outputs, the possibilities provided by the LLM can introduce a small bias in what experts perceive as ’what might have happened.’ Future research must recognize the importance of examining both the LLM’s biases in interpreting personal tracking data and the biases experts might develop when integrating those results.
@@ 9.4. Designing LLM agents that mimic experts’ sensemaking process
Although this study did not aim to enhance the accuracy of complex activity recognition and sensemaking, our insights into experts’ sensemaking processes may help inform the design of more effective AI-assisted systems for interpreting personal tracking data. Guiding LLMs with expert knowledge is a common approach for solving domain-specific problems, including techniques like adding instructions in prompts (ji2024hargpt, ; cosentino2024towards, ), creating expert agents (yang2024drhouse, ), and fine-tuning (kim2024health, ; cosentino2024towards, ). In the context of making sense of sensor data, HARGPT employs simple role-playing prompts without expert guidance to identify activities from raw IMU data (ji2024hargpt, ). Health-LLM evaluated zero/few-shot learning and fine-tuning with simple prompts on sensor data (kim2024health, ). Similarly, prior works have also explored how to encode sensor data to prompts: LLMSense and PhysioLLM use text-formatted sensor data and prompts to derive high-level inferences from sensor traces (ouyang2024llmsense, ; fang2024physiollm, ); Cosetino et al. incorporated domain knowledge into prompts and fine-tuned models using expert responses (cosentino2024towards, ). We build on previous works to provide an iterative and evolving approach to continuously include domain expertise as Data Interpretation Guidance in the prompt. While prior works have explored incorporating domain/expert knowledge, we observed that much of the expert knowledge is usually “static.”. From related work, it is clear that the complex sensemaking processe experts use when interpreting personal tracking data are under-documented and rarely considered to guide LLMs. Our study reveals that these sensemaking processes extend beyond basic knowledge, like the normal range of a person’s heart rate, and encompass the selection and integration of multiple data streams, handling data discrepancies and uncertainties, and strategies for information retrieval. These domain-specific insights can guide LLMs in various ways. For instance, understanding how experts initiate sensemaking – starting from changes rather than following a strict chronological order – could be used to optimize data retrieval for LLMs. Additionally, the differing definitions of anomalies between the LLM and experts highlight an additional area for improvement, involving extending the system to incorporate more task-specific and sophisticated anomaly detection algorithms. Our findings on how experts collaborate with LLMs and other informational modules to reach conclusions underscore how LLM detection models should be designed for real-world applications to support experts effectively. We hope that, although not the main focus of this study, our insights will help AI and LLM researchers recognize opportunities to enhance current detection systems and better support complex sensemaking tasks.
9.5. Replacing or assisting: Experts and AI collaboration in contextual reasoning of sensing data
Prior work using LLMs for contextual reasoning and sensor data understanding has predominantly focused on detection tasks – building AI systems that make inferences independently and compare their performance against human experts. The underlying goal in these cases is often to replace or reduce reliance on experts by demonstrating comparable or superior accuracy. In contrast, our work emphasizes a complementary vision where AI systems are designed to assist experts with contextual reasoning and understanding: a more interpretive and subjective process. We frame LLMs as collaborative partners that assist experts in synthesizing multimodal, complex, subjective, and context-dependent data, rather than simply predicting predefined outcomes. This distinction is critical, particularly given the inherent challenges in sensemaking such as ambiguity, trust, and cognitive bias. We recognize the importance of both goals and, in this section, we further discuss the different approaches taken in previous studies and our work to help guide future researchers in applying appropriate methods.
Previous works comparing LLM results with human input often focus on tasks with well-defined outcomes, such as activity recognition, stress scores, or diagnoses (yang2024drhouse, ; ji2024hargpt, ; kim2024health, ; cosentino2024towards, ; ouyang2024llmsense, ; fang2024physiollm, ). These tasks allow researchers to define clear metrics and conduct quantitative evaluations to claim success of certain techniques proposed in those studies, like prompt engineering techniques, multi-agent pipeline, fine-tuned model, and more. However, for more subjective tasks like summarizing a person’s day, there is no single “right” or “wrong” answer, thus making objective, quantitative comparisons difficult. For example, we used standardized NLP methods to measure the accuracy and consistency of the summaries, and found that expert-written summaries were significantly more diverse than those generated by LLM. However, these metrics are not enough to claim that LLM behaves better than experts. In real-world studies, both types of results are necessary for different applications. Comparative studies provide valuable insights for building fundamental models that support applications requiring automated outputs. For instance, activity recognition results generated by LLMs can be directly used as context for delivering personalized health interventions. Subjective results, on the other hand, are crucial for human-centered applications that involve multiple stakeholders, such as monitoring dashboards for researchers, daily summaries for self-reflection, or even clinical decision making. Adler et al., in their recent paper, proposed the concept of actionable sensing as an alternative to detection research, which brings passive sensing data on behavior and physiology into clinical encounters to improve patient care (adler2024beyond, ). We embrace a similar concept, where LLM’s extensive narrative capabilities create new opportunities for generating high-level content that is valuable for various stakeholders, extending beyond detection.
The techniques and applications of LLMs for the two goals are also different. In comparison studies, the focus is often on generating an accurate final result, whereas in AI aiming to assist experts, much of the value lies in providing explanations and building trust. As a result, other than approving the accuracy, ensuring alignment with experts’ mental models becomes a critical design implication for the LLM pipeline of these systems. In this paper, one of our key contributions is the experts’ sensemaking model of sensing data, where we suggest that future AI-assisted systems provide outputs that follow and support this process. We define several tasks where experts may need assistance in making sense of the data, and instruct LLMs to provide the desired information along with proper explanations. In works that focus on comparing LLM and expert outputs, experts’ sensemaking models or similar insights are often used as external knowledge resources to guide the LLM in making better decisions through prompt engineering, fine-tuning, and reinforcement learning.
To summarize, both goals of “replacing with” and “assisting” experts are valuable directions for future research; however, the latter, specifically related to enhancing understanding and contextual reasoning, has generally received less attention in the past. Recently, more and more researchers have recognized the significance of human-AI collaboration in understanding data, particularly within the healthcare context. For instance, recent work by Zhang et al. proposed a new clinician-AI collaboration paradigm instead of an existing human-AI competition paradigm to account for the high uncertainty and risks while diagnosing sepsis (zhang2024rethinking, ). We hope our work aligns with this direction and encourages more researchers to recognize the value and potential of AI in supporting experts in making sense of sensing data.
@@ 9.7. Limitations
The study has several limitations worth noting. While we achieved data saturation, our small sample size limits the generalizability of our findings, particularly as it relates to other health and behavioral outcomes. Although recent research suggests small sample sizes are acceptable in interpretive studies (crabtree2024h, ), future work should involve larger populations and diverse scenarios to validate the findings. While researchers reviewed all the LLM-generated content shown in the two user studies to ensure basic correctness without making any changes, we conducted a preliminary quantitative evaluation of the LLM results. A more thorough evaluation of LLM results and capabilities can be conducted in future research, which is beyond the scope of this study. Another limitation arises from the fact that experts spent more time on tasks when LLM content was included compared to when it was not. In our qualitative analysis, we found that this was partly due to experts treating LLM content seriously in their sensemaking process, which led them to spend more time on it. Additionally, although we provided a tutorial video at the beginning, there was still a learning curve in using the information effectively. A more longitudinal study is needed to assess the long-term usage of the prototype in real-world settings and determine whether it can ultimately save experts time in sensemaking tasks. Although we conducted A/B testing using a baseline system without LLM, we did not compare our system with other dashboards due to the specificity of the task. As a result, we presented the survey results as descriptive measurements and highlighted the positive qualitative feedback from experts, which aligned with the survey findings.